# Some Bayesian Perspectives on Clinical Trials

### v1

@@ 2.3 Computational advances: Pólya-Gamma augmentation
Polson et al. [2013] introduced the Pólya-Gamma data augmentation scheme, which renders the posterior for logistic regression conditionally Gaussian. The key identity is:
@@
For clinical trials with binary endpoints and moderate covariate adjustment, this scheme produces exact posterior samples with no tuning parameters, making real-time Bayesian monitoring computationally straightforward. The approach extends naturally to multinomial and negative binomial models.
@@ 3 Sequential Adaptive Designs
The frequentist approach to interim monitoring is group sequential testing. Pocock [1977] proposed equal critical values at each interim look; O’Brien and Fleming [1979] proposed conservative early boundaries that spend little alpha initially. Lan and DeMets [1983] unified these through the alpha-spending function $\alpha^{*}(t)$, which allocates the overall type-I error rate across information fractions $t\in[0,1]$. The spending function need not be specified until the interim analysis is conducted, giving the trialist flexibility in the timing and number of looks.
The Bayesian sequential framework differs in three respects. First, it conditions on the data actually observed rather than on a pre-specified schedule of looks, so inference is valid regardless of when or why the data were examined. Second, the stopping rule is based on predictive or posterior probabilities that answer the clinical question directly (“what is the probability that this treatment is superior?”) rather than on $p$-values adjusted for multiplicity. Third, the Bayesian framework extends naturally to response-adaptive randomization and multi-arm designs, whereas group sequential methods are largely restricted to fixed-allocation two-arm trials with pre-planned interim analyses.
@@ 3.1 Thompson sampling
Thompson [1933] proposed the earliest response-adaptive randomization rule. Given two treatments with unknown success probabilities $\tilde{p}_{1}$ and $\tilde{p}_{2}$, and prior data consisting of $r_{i}$ successes in $n_{i}$ patients for treatment $i$, Thompson computed
@@
Thompson’s motivation was explicitly ethical: “a considerable saving of individuals otherwise sacrificed to the inferior treatment might be effected.” Under uniform priors, the posterior for each treatment’s success rate is $\text{Beta}(r_{i}+1,s_{i}+1)$, and the allocation probabilities in (3) follow from the integral $\int_{0}^{1}\int_{0}^{\theta_{2}}f(\theta_{2}|\text{data})f(\theta_{1}|\text{data})\,d\theta_{1}\,d\theta_{2}$.
Thompson sampling was rediscovered by the technology industry in the 2000s and is now the dominant approach to online optimization problems—website design, advertisement placement, recommendation systems, and cloud service configuration [Scott, 2010]. Advances in Bayesian computation made randomized probability matching easy to implement for virtually any payoff distribution. If Thompson sampling has become the standard for allocating web traffic to competing designs, where the stakes are advertising revenue, it should a fortiori be the standard for allocating patients to competing treatments, where the stakes are human lives.
Figure 3 shows simulated Thompson sampling paths for a two-arm trial with control success rate $p_{0}=0.30$ and treatment success rate $p_{1}=0.45$. The allocation proportion drifts above the equal-randomization line as evidence accumulates, assigning more patients to the superior arm.
Figure 3: Thompson sampling in a two-arm trial ($p_{0}=0.30$, $p_{1}=0.45$). Each curve shows the running proportion of patients assigned to the treatment arm across eight independent simulations. Under equal randomization (dashed line), half of all patients receive the inferior control. Thompson sampling progressively shifts allocation toward the superior arm as evidence accumulates.
The Thall-Wathen generalization raises the posterior probabilities to a power $\tau$ before normalizing:
@@
where $p_{k,n}=\Pr(\pi_{k}>\pi_{j}\text{ for all }j\neq k\mid D_{n})$ is the posterior probability that arm $k$ is best after $n$ patients, and $K$ is the number of arms. When $\tau=0$, this reduces to equal randomization; when $\tau\to\infty$, it assigns all patients to the current leader. The tuning parameter $\tau$ controls the exploration-exploitation tradeoff inherent in the multi-armed bandit formulation of clinical trials.
@@ 3.2 Predictive probabilities and stopping rules
A trial’s most consequential decision is when to stop. Bayesian predictive probabilities answer the question trial teams actually face: given current data, what is the probability that the trial will reach a definitive conclusion if continued? The idea has an early precedent in bioequivalence testing: Racine-Poon et al. [1987] proposed a two-stage procedure in which first-stage data are used to compute the predictive probability of establishing bioequivalence in a second stage, and the second-stage sample size $n_{2}$ is chosen—or the study abandoned—accordingly. Their simulations showed that the two-stage design roughly doubles the success rate of single-stage studies at a comparable average sample size.
Let $Z_{n}$ denote the test statistic at the interim analysis with information fraction $r=I_{n}/I_{N}$, where $I_{n}$ is the current information and $I_{N}$ is the information at the planned final analysis. Marion et al. [2025] show that the predictive probability of trial success can be approximated as
@@
where $p_{n}$ is the current $p$-value (or $1-\mathcal{P}_{n}$ for Bayesian posterior probability $\mathcal{P}_{n}$) and $\alpha$ is the significance threshold. This approximation is exact when the test statistic is Gaussian with known information fraction and a flat prior, and accurate across a range of endpoint types—continuous, binary, time-to-event, and ordinal—as verified by simulation against full Monte Carlo imputation [Marion et al., 2025].
@@ 3.3 Backward induction on sufficient statistics
Christen and Nakamura [2003], building on Berry et al. [2000] and Carlin et al. [1998], identified a key structural insight: when the utility function depends on data only through a sufficient statistic, the exponentially growing decision tree collapses to a polynomial-size table.
@@
Carlin et al. [1998] applied this sufficient-statistic reduction to clinical trials directly, developing sequential decision-theoretic methods for treatment allocation. Their approach demonstrated that backward induction—the “theoretically sound and generally accepted procedure for Bayesian sequential analysis” that is “hardly ever implemented due to its great complexities” [Berger, 1985]—becomes feasible when the problem structure is exploited.
@@ 3.4 The Normal model and gridding: a worked example
We develop the backward induction in full for the clinical trial problem of Brockwell and Kadane [2003], following Carlin et al. [1998]. This example makes the sufficient-statistic reduction concrete and illustrates the gridding algorithm that makes optimal sequential design computationally tractable.
@@
with $\sigma^{2}$ known. The decision at each stage is either to stop and choose a terminal action $d\in\{d_{0},d_{1}\}$ (declare for control or treatment) or to continue sampling at cost $c$ per patient. The loss for terminal decision $d_{1}$ (favor treatment) when the true effect is $\theta$ is $R(\theta,d_{1})$, and similarly for $d_{0}$; a natural choice is $R(\theta,d_{1})=-\theta$ (reward equals effect size if treatment is chosen) and $R(\theta,d_{0})=0$.
@@
Brockwell and Kadane [2003] evaluate (14) on a fine grid $\mathcal{G}=\{s_{1},\ldots,s_{G}\}$ over the range of $S_{n}$. In their clinical trial example, $\mathcal{G}$ spans [6,6][-6,6] with G=12, [^0] G=12{,}000 points. The algorithm proceeds backward from $T$:
@@
    by numerical quadrature over the Gaussian transition (12). Set $V_{n}(s_{j})=\min\{h_{n}(s_{j}),Q_{n}(s_{j})\}$.
@@
The optimal policy partitions the $(n,S_{n})$ plane into three regions: stop and choose treatment ($S_{n}$ large), stop and choose control ($S_{n}$ small), and continue (intermediate $S_{n}$). These regions form a triangle that narrows with $n$: early in the trial, the continuation region is wide (much can still be learned); as $n$ approaches $T$, the boundaries converge and the trial is forced to a decision. Figure 4 shows the result for the Normal model with $\sigma^{2}=\sigma_{0}^{2}=1$ and sampling cost $c=0.02$, together with six simulated trial paths that terminate when they cross a boundary. This is the Bayesian analog of the triangular continuation region in frequentist group sequential designs [Pocock, 1977], but here the boundaries are derived from the loss function and the prior rather than from alpha-spending constraints.
Figure 4: Optimal stopping regions from backward induction on the Normal model (9) with $\sigma^{2}=\sigma_{0}^{2}=1$ and sampling cost $c=0.02$ per patient. The upper region (green) indicates stopping in favor of treatment; the lower region (red) indicates stopping in favor of control; the middle region indicates continuing. Colored curves show six simulated trial paths with different true $\theta$ values, each terminating when it crosses a boundary.

### v2

@@ 2.3 Pólya-Gamma augmentation
Polson et al. (2013) introduced the Pólya-Gamma data augmentation scheme, a special case of the variance-mean mixture framework for non-Gaussian regression (Polson and Scott, 2013), which renders the posterior for logistic regression conditionally Gaussian. The key identity is:
@@
For clinical trials with binary endpoints and moderate covariate adjustment, this scheme produces posterior samples via a Gibbs sampler with no tuning parameters, making real-time Bayesian monitoring computationally straightforward. The approach extends naturally to multinomial and negative binomial models.
Multi-centre trials.
A direct application arises in multi-centre binary-response studies, where the same treatment comparison is conducted across $N$ centres with potentially different response rates. Let $y_{ij}$ denote the number of successes among $n_{ij}$ patients at centre $i$ on arm $j\in\{1,2\}$ (treatment, control), with log-odds $\psi_{ij}=\log\{p_{ij}/(1-p_{ij})\}$. The hierarchical logistic-normal model places a bivariate normal prior on each centre’s log-odds vector:
$\psi_{i}=(\psi_{i1},\psi_{i2})^{\prime}\mid\mu,\Sigma\sim N(\mu,\Sigma),\qquad p(\mu)\propto 1,\qquad\Sigma\sim\text{IW}(d,B),$ (6)
where $\mu=(\mu_{1},\mu_{2})^{\prime}$ represents the overall treatment and control log-odds, the improper flat prior on $\mu$ is standard, and $\Sigma$ captures between-centre variability. Skene and Wakefield (1990) proposed this model and analysed it via numerical integration; Polson and Scott (2011) showed that Pólya-Gamma augmentation renders the posterior conditionally conjugate, yielding a tuning-free Gibbs sampler.
Applying the Pólya-Gamma identity (5) to each binomial likelihood factor $p(y_{ij}\mid\psi_{ij})\propto e^{y_{ij}\psi_{ij}}/(1+e^{\psi_{ij}})^{n_{ij}}$ introduces latent variables $\omega_{ij}\mid\psi_{ij}\sim\text{PG}(n_{ij},\psi_{ij})$. Conditional on $\boldsymbol{\omega}$, the centre-level log-odds have a Gaussian full conditional:
$\psi_{i}\mid\boldsymbol{\omega}_{i},\mu,\Sigma,D\sim N(\mathbf{m}_{i},\mathbf{V}_{i}),\qquad\mathbf{V}_{i}^{-1}=\boldsymbol{\Omega}_{i}+\Sigma^{-1},\quad\mathbf{m}_{i}=\mathbf{V}_{i}(\boldsymbol{\kappa}_{i}+\Sigma^{-1}\mu),$ (7)
where $\boldsymbol{\Omega}_{i}=\text{diag}(\omega_{i1},\omega_{i2})$ and $\boldsymbol{\kappa}_{i}=(y_{i1}-n_{i1}/2,\;y_{i2}-n_{i2}/2)^{\prime}$. The hyperparameters update conjugately: $\mu\mid\Psi,\Sigma\sim N(\bar{\psi},\Sigma/N)$ and $\Sigma\mid\Psi,\mu\sim\text{IW}(d+N,B+\sum_{i=1}^{N}(\psi_{i}-\mu)(\psi_{i}-\mu)^{\prime})$.
The structure of (7) also clarifies the connection to the conditional means prior (4). If one replaces the logistic-normal prior in (6) by independent $\text{Beta}(a_{ij},b_{ij})$ priors on the success probabilities $p_{ij}$—equivalently, independent Fisher $Z(a_{ij},b_{ij})$ priors on the log-odds $\psi_{ij}$—then Pólya-Gamma augmentation absorbs these priors into the Gaussian conditional as $a_{ij}+b_{ij}$ pseudo-observations (Polson and Scott, 2011). The elicited Beta hyperparameters at design points become pseudo-data that enter the Gibbs sampler on the same footing as observed counts, exactly as in the Bedrick–Christensen–Johnson framework. This provides a computationally trivial route from the conditional means prior specification of Section 2 to full posterior inference, whether the hierarchical coupling across centres is through a shared logistic-normal prior on $\psi$ or through independent Z-priors with common hyperparameters.
Figure 3: Joint posterior for $\mu=(\mu_{1},\mu_{2})^{\prime}$ in the hierarchical logistic-normal model (6) fitted to the eight-centre topical cream data of Skene and Wakefield (1990) via Pólya-Gamma Gibbs sampling. Each grey dot is a posterior draw; the diagonal line marks $\mu_{1}=\mu_{2}$ (equal treatment and control). The mass below the diagonal ($\Pr(\mu_{1}>\mu_{2}\mid\text{data})=0.985$) indicates strong evidence of treatment efficacy despite small per-centre samples.
Figure 3 illustrates the general small-sample advantage on data from an eight-centre study of topical cream effectiveness (Skene and Wakefield, 1990). Individual centres are small—5 to 37 patients per arm—and most are individually inconclusive (although one or two reach nominal significance by Fisher’s exact test, none survive a Bonferroni correction for eight comparisons). Yet the hierarchical model borrows strength across centres, and the posterior for $\mu$ places 98.5% of its mass in the region $\mu_{1}>\mu_{2}$ (below the diagonal in the figure), reaching a strong pooled conclusion that no single centre could support on its own. The Bayesian shrinkage is especially important for centres 5 and 6, where zero control-arm successes would produce infinite maximum-likelihood log-odds estimates; the hierarchical prior regularises these towards the population mean. The mechanism is visible in the precision matrix of (7): each centre’s posterior precision $\mathbf{V}_{i}^{-1}=\boldsymbol{\Omega}_{i}+\Sigma^{-1}$ combines local data (through $\boldsymbol{\Omega}_{i}$, which grows with sample size) with information from all other centres (through $\Sigma^{-1}$, which acts as an informative prior). A centre with few patients has small $\boldsymbol{\Omega}_{i}$ and is therefore pulled strongly towards the population mean $\mu$; a centre with many patients retains its own estimate. This is why the Bayesian framework reaches conclusions quickly in either direction: centres with clear evidence for or against the treatment contribute that evidence to the population-level inference, while centres with ambiguous data borrow from the collective. With an informative prior on $\mu$ calibrated to historical response rates—as in the Bedrick framework of Section 2—the borrowing starts before the first patient is enrolled, enabling each centre to reach a decision with even fewer observations.
@@ 3 Sequential Adaptive Designs
The frequentist approach to interim monitoring is group sequential testing. Pocock (1977) proposed equal critical values at each interim look; O’Brien and Fleming (1979) proposed conservative early boundaries that spend little alpha initially. Lan and DeMets (1983) unified these through the alpha-spending function $\alpha^{*}(t)$, which allocates the overall type-I error rate across information fractions $t\in[0,1]$. The spending function need not be specified until the interim analysis is conducted, giving the trialist flexibility in the timing and number of looks.
The Bayesian sequential framework differs in three respects. First, it conditions on the data actually observed rather than on a pre-specified schedule of looks, so inference under the assumed model is valid regardless of when or why the data were examined. Second, the stopping rule is based on predictive or posterior probabilities that answer the clinical question directly (“what is the probability that this treatment is superior?”) rather than on $p$-values adjusted for multiplicity. Third, the Bayesian framework extends naturally to response-adaptive randomisation and multi-arm designs; frequentist group sequential methods can accommodate multiple arms and adaptive allocation, but doing so requires increasingly elaborate multiplicity adjustments, whereas the Bayesian approach handles these extensions through the posterior, though confirmatory trials still require simulation-based calibration of operating characteristics including type I error.
@@ 3.1 Thompson sampling
Thompson (1933) proposed the earliest response-adaptive randomisation rule. Given two treatments with unknown success probabilities $\tilde{p}_{1}$ and $\tilde{p}_{2}$, and prior data consisting of $r_{i}$ successes in $n_{i}$ patients for treatment $i$, Thompson computed
@@
Thompson’s motivation was explicitly ethical: “a considerable saving of individuals otherwise sacrificed to the inferior treatment might be effected.” Under uniform priors, the posterior for each treatment’s success rate is $\text{Beta}(r_{i}+1,s_{i}+1)$, and the allocation probabilities in (8) follow from the integral $\int_{0}^{1}\int_{0}^{\theta_{2}}\pi(\theta_{2}|\text{data})\pi(\theta_{1}|\text{data})\,d\theta_{1}\,d\theta_{2}$.
Thompson sampling was rediscovered by the technology industry in the 2000s and is now widely adopted for online optimisation problems (Scott, 2010). Advances in Bayesian computation made randomised probability matching easy to implement for virtually any payoff distribution. The adoption of Thompson sampling for allocating web traffic—where the stakes are advertising revenue—suggests that the case for adaptive allocation is at least as strong in clinical trials, where the stakes are patient outcomes.
Figure 4 shows simulated Thompson sampling paths for a two-arm trial with control success rate $p_{0}=0.30$ and treatment success rate $p_{1}=0.45$. The allocation proportion drifts above the equal-randomisation line as evidence accumulates, assigning more patients to the superior arm.
Figure 4: Thompson sampling in a two-arm trial ($p_{0}=0.30$, $p_{1}=0.45$). Each curve shows the running proportion of patients assigned to the treatment arm across eight independent simulations. Under equal randomisation (dashed line), half of all patients receive the inferior control. Thompson sampling progressively shifts allocation toward the superior arm as evidence accumulates.
The Thall and Wathen (2007) generalisation raises the posterior probabilities to a power $\tau$ before normalising:
@@
where $p_{k,n}=\Pr(\pi_{k}>\pi_{j}\text{ for all }j\neq k\mid D_{n})$ is the posterior probability that arm $k$ is best after $n$ patients, and $K$ is the number of arms. When $\tau=0$, this reduces to equal randomisation; when $\tau\to\infty$, it assigns all patients to the current leader. The tuning parameter $\tau$ controls the exploration-exploitation trade-off inherent in the multi-armed bandit formulation of clinical trials.
@@ 3.2 Predictive probabilities and stopping rules
A trial’s most consequential decision is when to stop. Bayesian predictive probabilities answer the question trial teams actually face: given current data, what is the probability that the trial will reach a definitive conclusion if continued? The idea has an early precedent in bioequivalence testing: Racine-Poon et al. (1987) proposed a two-stage procedure in which first-stage data are used to compute the predictive probability of establishing bioequivalence in a second stage, and the second-stage sample size $n_{2}$ is chosen—or the study abandoned—accordingly. Their simulations showed that the two-stage design roughly doubles the success rate of single-stage studies at a comparable average sample size.
Let $Z_{n}$ denote the test statistic at the interim analysis with information fraction $r=I_{n}/I_{N}$, where $I_{n}$ is the current information and $I_{N}$ is the information at the planned final analysis. Marion et al. (2025) show that the predictive probability of trial success can be approximated as
@@
where $p_{n}$ is the one-sided $p$-value at the interim (equivalently, $p_{n}=1-\mathcal{P}_{n}$ when $\mathcal{P}_{n}$ is the Bayesian posterior probability of superiority under a flat prior), $r=I_{n}/I_{N}$ is the observed information fraction, and $\alpha$ is the one-sided significance threshold for the final analysis. The approximation is exact when the test statistic follows a Gaussian information-monitoring process with known $r$ and a flat prior, and accurate across a range of endpoint types—binary, time-to-event, and ordinal—as verified by simulation against full Monte Carlo imputation (Marion et al., 2025).
@@ 3.3 Backward induction on sufficient statistics
Christen and Nakamura (2003), building on Berry et al. (2000) and Carlin et al. (1998), identified a key structural insight: when the utility function depends on data only through a sufficient statistic, the exponentially growing decision tree collapses to a polynomial-size table.
@@
Carlin et al. (1998) applied this sufficient-statistic reduction to clinical trials directly, developing sequential decision-theoretic methods for treatment allocation. Their approach demonstrated that backward induction—the “theoretically sound and generally accepted procedure for Bayesian sequential analysis” that is “hardly ever implemented due to its great complexities” (Berger, 1985)—becomes feasible when the problem structure is exploited.
@@ 3.4 The Normal model and gridding: a worked example
We develop the backward induction in full for the clinical trial problem of Brockwell and Kadane (2003), following Carlin et al. (1998). This example makes the sufficient-statistic reduction concrete and illustrates the gridding algorithm that makes optimal sequential design computationally tractable.
@@
with $\sigma^{2}$ known. The decision at each stage is either to stop and choose a terminal action $d\in\{d_{0},d_{1}\}$ (declare for control or treatment) or to continue sampling at cost $c$ per patient. The loss for terminal decision $d_{1}$ (favour treatment) when the true effect is $\theta$ is $R(\theta,d_{1})$, and similarly for $d_{0}$; a natural choice is $R(\theta,d_{1})=-\theta$ (reward equals effect size if treatment is chosen) and $R(\theta,d_{0})=0$.
@@
Brockwell and Kadane (2003) evaluate (19) on a fine grid $\mathcal{G}=\{s_{1},\ldots,s_{G}\}$ over the range of $S_{n}$. In their clinical trial example, $\mathcal{G}$ spans [6,6][-6,6] with G=12, [^0] G=12{,}000 points. The algorithm proceeds backward from $T$:
@@
    by numerical quadrature over the Gaussian transition (17). Set
    $V_{n}(s_{j})=\min\{h_{n}(s_{j}),Q_{n}(s_{j})\}.$
@@
The optimal policy partitions the $(n,S_{n})$ plane into three regions: stop and choose treatment ($S_{n}$ large), stop and choose control ($S_{n}$ small), and continue (intermediate $S_{n}$). These regions form a triangle that narrows with $n$: early in the trial, the continuation region is wide (much can still be learned); as $n$ approaches $T$, the boundaries converge and the trial is forced to a decision. Figure 5 shows the result for the Normal model with $\sigma^{2}=4$, $\sigma_{0}^{2}=1$, and sampling cost $c=0.005$, together with eight simulated trial paths that terminate when they cross a boundary. The shape resembles the triangular continuation region in frequentist group sequential designs, but here the boundaries are derived from the loss function and the prior rather than from alpha-spending constraints.
Figure 5: Optimal stopping regions from backward induction on the Normal model (14) with $\sigma^{2}=4$, $\sigma_{0}^{2}=1$, and sampling cost $c=0.005$ per patient. The upper region (green) indicates stopping in favour of treatment; the lower region (red) indicates stopping in favour of control; the middle region indicates continuing. Coloured curves show eight simulated trial paths with different true $\theta$ values, each terminating when it crosses a boundary.